\begin{thebibliography}{55}
\providecommand{\natexlab}[1]{#1}

\bibitem[{Amodei(2024)}]{amodei2024machines}
Dario Amodei. 2024.
\newblock \href {https://www.darioamodei.com/essay/machines-of-loving-grace} {Machines of loving grace: How {AI} could transform the world for the better}.
\newblock Online essay.
\newblock \url{https://www.darioamodei.com/essay/machines-of-loving-grace}.

\bibitem[{Bender and Koller(2020)}]{bender-koller-2020-climbing}
Emily~M. Bender and Alexander Koller. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.463} {Climbing towards {NLU}: {On} meaning, form, and understanding in the age of data}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 5185--5198, Online. Association for Computational Linguistics.

\bibitem[{Blili-Hamelin et~al.(2025)Blili-Hamelin, Graziul, Hancox-Li, Hazan, El-Mhamdi, Ghosh, Heller, Metcalf, Murai, Salvaggio, Smart, Snider, Tighanimine, Ringer, Mitchell, and Dori-Hacohen}]{blilihamelin2025stoptreatingaginorthstar}
Borhane Blili-Hamelin, Christopher Graziul, Leif Hancox-Li, Hananel Hazan, El-Mahdi El-Mhamdi, Avijit Ghosh, Katherine Heller, Jacob Metcalf, Fabricio Murai, Eryk Salvaggio, Andrew Smart, Todd Snider, Mariame Tighanimine, Talia Ringer, Margaret Mitchell, and Shiri Dori-Hacohen. 2025.
\newblock \href {https://arxiv.org/abs/2502.03689} {Stop treating `{AGI}' as the north-star goal of ai research}.
\newblock \emph{Preprint}, arXiv:2502.03689.

\bibitem[{Blodgett et~al.(2020)Blodgett, Barocas, Daum{\'e}~III, and Wallach}]{blodgett-etal-2020-language}
Su~Lin Blodgett, Solon Barocas, Hal Daum{\'e}~III, and Hanna Wallach. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.acl-main.485} {Language (technology) is power: A critical survey of ``bias'' in {NLP}}.
\newblock In \emph{Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics}, pages 5454--5476, Online. Association for Computational Linguistics.

\bibitem[{Boyer(1996)}]{boyer1996makes}
Pascal Boyer. 1996.
\newblock What makes anthropomorphism natural: Intuitive ontology and cultural representations.
\newblock \emph{Journal of the Royal Anthropological Institute}, pages 83--97.

\bibitem[{Buchanan and Smith(1988)}]{buchanan1988fundamentals}
Bruce~G Buchanan and Reid~G Smith. 1988.
\newblock Fundamentals of expert systems.
\newblock \emph{Annual Review of Computer Science}, 3(1):23--58.

\bibitem[{Clarke et~al.(2023)Clarke, Milmo, and Blight}]{clarke2023_how_ai_chatbots_work}
Seán Clarke, Dan Milmo, and Garry Blight. 2023.
\newblock \href {https://www.theguardian.com/technology/ng-interactive/2023/nov/01/how-ai-chatbots-like-chatgpt-or-bard-work-visual-explainer} {How {AI} chatbots like {ChatGPT} or {Bard} work – visual explainer}.
\newblock Accessed: 2025-06-19.

\bibitem[{Coghlan et~al.(2023)Coghlan, Leins, Sheldrick, Cheong, Gooding, and D'Alfonso}]{coghlan2023chat}
Simon Coghlan, Kobi Leins, Susie Sheldrick, Marc Cheong, Piers Gooding, and Simon D'Alfonso. 2023.
\newblock To chat or bot to chat: Ethical issues with using chatbots in mental health.
\newblock \emph{Digital health}, 9:20552076231183542.

\bibitem[{Cormack et~al.(2008)}]{cormack2008email}
Gordon~V Cormack and 1 others. 2008.
\newblock Email spam filtering: A systematic review.
\newblock \emph{Foundations and Trends{\textregistered} in Information Retrieval}, 1(4):335--455.

\bibitem[{Crevier(1993)}]{crevier1993ai}
Daniel Crevier. 1993.
\newblock \emph{AI: The Tumultuous History of the Search for Artificial Intelligence}, 1st edition.
\newblock Basic Books, New York, NY.

\bibitem[{Edwards and Shafer(2022)}]{edwards2022lamps}
Alyssa~D Edwards and Daniel~M Shafer. 2022.
\newblock When lamps have feelings: Empathy and anthropomorphism toward inanimate objects in animated films.
\newblock \emph{Projections}, 16(2):27--52.

\bibitem[{Faverio and Tyson(2023)}]{FaverioTyson2023}
Michelle Faverio and Alec Tyson. 2023.
\newblock \href {https://www.pewresearch.org/short-reads/2023/11/21/what-the-data-says-about-americans-views-of-artificial-intelligence/} {What the data says about americans’ views of artificial intelligence}.
\newblock Pew Research Center – Short Reads.

\bibitem[{Fontaine et~al.(2019)Fontaine, Maheu-Cadotte, Lavall{\'e}e, Mailhot, Rouleau, Bouix-Picasso, and Bourbonnais}]{fontaine2019communicating}
Guillaume Fontaine, Marc-Andr{\'e} Maheu-Cadotte, Andr{\'e}ane Lavall{\'e}e, Tanya Mailhot, Genevi{\`e}ve Rouleau, Julien Bouix-Picasso, and Anne Bourbonnais. 2019.
\newblock \href {https://doi.org/10.2196/14447} {Communicating science in the digital and social media ecosystem: Scoping review and typology of strategies used by health scientists}.
\newblock \emph{JMIR Public Health and Surveillance}, 5(3):e14447.

\bibitem[{Foote(2021)}]{foote2021briefml}
Keith~D. Foote. 2021.
\newblock \href {https://www.dataversity.net/a-brief-history-of-machine-learning/} {A brief history of machine learning}.
\newblock Online article.
\newblock \url{https://www.dataversity.net/a-brief-history-of-machine-learning/}.

\bibitem[{Fried(2025)}]{Fried2025}
Ina Fried. 2025.
\newblock \href {https://www.axios.com/2025/03/06/exclusive-russian-disinfo-floods-ai-chatbots-study-finds} {Exclusive: Russian disinformation floods {AI} chatbots, study finds}.
\newblock \emph{Axios}.
\newblock Based on a NewsGuard report.

\bibitem[{Ghosh et~al.(2024)Ghosh, Venkit, Gautam, Wilson, and Caliskan}]{Ghosh2024}
Sourojit Ghosh, Pranav~Narayanan Venkit, Sanjana Gautam, Shomir Wilson, and Aylin Caliskan. 2024.
\newblock \href {https://doi.org/10.1609/aies.v7i1.31651} {Do generative {AI} models output harm while representing non-western cultures: Evidence from a community‑centered approach}.
\newblock In \emph{Proceedings of the AAAI/ACM Conference on AI, Ethics, and Society (AIES)}, volume~7 of \emph{AIES '24}, pages 476--489.

\bibitem[{Govindan and Shivaprasad(1990)}]{GOVINDAN1990671}
V.K Govindan and A.P Shivaprasad. 1990.
\newblock \href {https://doi.org/10.1016/0031-3203(90)90091-X} {Character recognition — a review}.
\newblock \emph{Pattern Recognition}, 23(7):671--683.

\bibitem[{Grozdanoff(2023)}]{grozdanoff2023looming}
Boris~D Grozdanoff. 2023.
\newblock The looming shadow: Taking seriously potential existential threats brought about by artificial intelligence.
\newblock \emph{Ethical Studies}, pages 100--106.

\bibitem[{Hayes(2025)}]{hayes2025_ai_jobs}
Adam Hayes. 2025.
\newblock \href {https://www.investopedia.com/could-ai-be-coming-for-your-job-11749570} {Is {AI} coming for your job? here's how to tell}.
\newblock \emph{Investopedia}.
\newblock Accessed: 2025-06-19.

\bibitem[{Heikkilä(2025)}]{heikkila2025_mitchell_agi_snake_oil}
Melissa Heikkilä. 2025.
\newblock Margaret mitchell: artificial general intelligence is ‘just vibes and snake oil’.
\newblock \emph{Financial Times}.
\newblock Accessed: 2025-06-19.

\bibitem[{Hendler(2008)}]{4475849}
James Hendler. 2008.
\newblock \href {https://doi.org/10.1109/MIS.2008.20} {Avoiding another {AI} winter}.
\newblock \emph{IEEE Intelligent Systems}, 23(2):2--4.

\bibitem[{Hill and Hengchen(2019)}]{hill2019quantifying}
Mark~J Hill and Simon Hengchen. 2019.
\newblock Quantifying the impact of dirty ocr on historical text analysis: Eighteenth century collections online as a case study.
\newblock \emph{Digital Scholarship in the Humanities}, 34(4):825--843.

\bibitem[{Horowitch(2025)}]{horowitch2025computersciencebubble}
Rose Horowitch. 2025.
\newblock \href {https://www.theatlantic.com/economy/archive/2025/06/computer-science-bubble-ai/683242/} {The computer‑science bubble is bursting}.
\newblock \emph{The Atlantic}.

\bibitem[{Huang et~al.(2025)Huang, Yu, Ma, Zhong, Feng, Wang, Chen, Peng, Feng, Qin, and Liu}]{huang2025survey}
Lei Huang, Weijiang Yu, Weitao Ma, Weihong Zhong, Zhangyin Feng, Haotian Wang, Qianglong Chen, Weihua Peng, Xiaocheng Feng, Bing Qin, and Ting Liu. 2025.
\newblock \href {https://doi.org/10.1145/3703155} {A survey on hallucination in large language models: Principles, taxonomy, challenges, and open questions}.
\newblock \emph{ACM Trans. Inf. Syst.}, 43(2).

\bibitem[{Hudson and Franklin(2023)}]{hudson2023}
Simon Hudson and Matija Franklin. 2023.
\newblock \href {https://arxiv.org/abs/2308.16377} {Science communications for explainable artificial intelligence}.
\newblock \emph{Preprint}, arXiv:2308.16377.

\bibitem[{James et~al.(1973)James, James, Stuart, Roger, and Christopher}]{james1973artificial}
Lighthill James, Lighthill James, Sutherland Stuart, Needham Roger, and Longuet-Higgins Christopher. 1973.
\newblock Artificial intelligence: A general survey.
\newblock \emph{Science Research Council}.

\bibitem[{Kosmyna et~al.(2025)Kosmyna, Hauptmann, Yuan, Situ, Liao, Beresnitzky, Braunstein, and Maes}]{kosmyna2025brainchatgptaccumulationcognitive}
Nataliya Kosmyna, Eugene Hauptmann, Ye~Tong Yuan, Jessica Situ, Xian-Hao Liao, Ashly~Vivian Beresnitzky, Iris Braunstein, and Pattie Maes. 2025.
\newblock \href {https://arxiv.org/abs/2506.08872} {Your brain on {ChatGPT}: Accumulation of cognitive debt when using an ai assistant for essay writing task}.
\newblock \emph{Preprint}, arXiv:2506.08872.

\bibitem[{Kuncel and Rigdon(2013)}]{kuncel2013communicating}
Nathan~R. Kuncel and J.~Rigdon. 2013.
\newblock Communicating research findings.
\newblock In Neal~W. Schmitt, Scott Highhouse, and Irving~B. Weiner, editors, \emph{Handbook of Psychology: Industrial and Organizational Psychology}, 2nd edition, volume~12, pages 43--58. John Wiley \& Sons, Inc., Hoboken, NJ.

\bibitem[{Liddy et~al.(1994)Liddy, Paik, and Yu}]{10.1145/183422.183425}
Elizabeth~D. Liddy, Woojin Paik, and Edmund~S. Yu. 1994.
\newblock \href {https://doi.org/10.1145/183422.183425} {Text categorization for multiple users based on semantic features from a machine-readable dictionary}.
\newblock \emph{ACM Trans. Inf. Syst.}, 12(3):278–295.

\bibitem[{Marois and Ivanoff(2005)}]{marois2005capacity}
Ren{\'e} Marois and Jason Ivanoff. 2005.
\newblock Capacity limits of information processing in the brain.
\newblock \emph{Trends in cognitive sciences}, 9(6):296--305.

\bibitem[{Mitchell and Krakauer(2023)}]{mitchell2023debate}
Melanie Mitchell and David~C Krakauer. 2023.
\newblock The debate over understanding in {AI's} large language models.
\newblock \emph{Proceedings of the National Academy of Sciences}, 120(13):e2215907120.

\bibitem[{Mosbach et~al.(2024)Mosbach, Gautam, Vergara~Browne, Klakow, and Geva}]{mosbach-etal-2024-insights}
Marius Mosbach, Vagrant Gautam, Tom{\'a}s Vergara~Browne, Dietrich Klakow, and Mor Geva. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.emnlp-main.181} {From insights to actions: The impact of interpretability and analysis research on {NLP}}.
\newblock In \emph{Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing}, pages 3078--3105, Miami, Florida, USA. Association for Computational Linguistics.

\bibitem[{Mueller(2024)}]{mueller2024myth}
Milton Mueller. 2024.
\newblock The myth of {AGI}.
\newblock \emph{Internet Governance Project}.

\bibitem[{Narayanan~Venkit et~al.(2024)Narayanan~Venkit, Chakravorti, Gupta, Biggs, Srinath, Goswami, Rajtmajer, and Wilson}]{narayanan-venkit-etal-2024-audit}
Pranav Narayanan~Venkit, Tatiana Chakravorti, Vipul Gupta, Heidi Biggs, Mukund Srinath, Koustava Goswami, Sarah Rajtmajer, and Shomir Wilson. 2024.
\newblock \href {https://doi.org/10.18653/v1/2024.emnlp-main.375} {An audit on the perspectives and challenges of hallucinations in {NLP}}.
\newblock In \emph{Proceedings of the 2024 Conference on Empirical Methods in Natural Language Processing}, pages 6528--6548, Miami, Florida, USA. Association for Computational Linguistics.

\bibitem[{{National Academies of Sciences, Engineering, and Medicine}(2017)}]{national2017communicating}
{National Academies of Sciences, Engineering, and Medicine}. 2017.
\newblock \href {https://doi.org/10.17226/23674} {Communicating science effectively: A research agenda}.

\bibitem[{Novelli et~al.(2024)Novelli, Casolari, Hacker, Spedicato, and Floridi}]{novelli2024generative}
Claudio Novelli, Federico Casolari, Philipp Hacker, Giorgio Spedicato, and Luciano Floridi. 2024.
\newblock Generative {AI} in {EU} law: Liability, privacy, intellectual property, and cybersecurity.
\newblock \emph{Computer Law \& Security Review}, 55:106066.

\bibitem[{O'Grady(2025)}]{OGrady2025_unethical_AI_Reddit}
Cathleen O'Grady. 2025.
\newblock \href {https://doi.org/10.1126/science.ady8074} {`{U}nethical' {AI} research on reddit under fire}.
\newblock \emph{Science}, 388(6747):570--571.
\newblock Perspective.

\bibitem[{Osawa et~al.(2022)Osawa, Miyamoto, Hase, Saijo, Fukuchi, and Miyake}]{osawa2022visions}
Hirotaka Osawa, Dohjin Miyamoto, Satoshi Hase, Reina Saijo, Kentaro Fukuchi, and Yoichiro Miyake. 2022.
\newblock Visions of artificial intelligence and robots in science fiction: a computational analysis.
\newblock \emph{International Journal of Social Robotics}, 14(10):2123--2133.

\bibitem[{Piquard(2024)}]{piquard2024_bender_parrots_lemonde}
Alexandre Piquard. 2024.
\newblock Chatbots are like parrots, they repeat without understanding.
\newblock \emph{Le Monde (English edition)}.
\newblock Interview with Emily M. Bender; Accessed: 2025-06-19.

\bibitem[{Press(2025)}]{press2025peakai}
Gil Press. 2025.
\newblock \href {https://www.forbes.com/sites/gilpress/2025/03/30/are-we-at-peak-ai-bubble-and-the-cusp-of-ai-moment/} {Are we at peak {AI} bubble and the cusp of {AI} moment?}
\newblock \emph{Forbes}.

\bibitem[{Riloff and Lehnert(1992)}]{riloff1992classifying}
Ellen Riloff and Wendy~G Lehnert. 1992.
\newblock Classifying texts using relevancy signatures.
\newblock In \emph{AAAI}, pages 329--334.

\bibitem[{Rogers and Augenstein(2020)}]{rogers-augenstein-2020-improve}
Anna Rogers and Isabelle Augenstein. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.findings-emnlp.112} {What can we do to improve peer review in {NLP}?}
\newblock In \emph{Findings of the Association for Computational Linguistics: EMNLP 2020}, pages 1256--1262, Online. Association for Computational Linguistics.

\bibitem[{Rogers(2025)}]{rogers2025_ai_bias_multilingual}
Reece Rogers. 2025.
\newblock Ai is spreading old stereotypes to new languages and cultures.
\newblock \emph{WIRED}.
\newblock Accessed: 2025-06-19.

\bibitem[{Roose(2023)}]{roose2023does}
Kevin Roose. 2023.
\newblock How does {ChatGPT} really work.
\newblock \emph{New York Times}, 28.

\bibitem[{Samuel(1959)}]{10.1147/rd.33.0210}
A.~L. Samuel. 1959.
\newblock \href {https://doi.org/10.1147/rd.33.0210} {Some studies in machine learning using the game of checkers}.
\newblock \emph{IBM J. Res. Dev.}, 3(3):210–229.

\bibitem[{Schwartz(1998)}]{schwartz1998web}
Candy Schwartz. 1998.
\newblock Web search engines.
\newblock \emph{Journal of the American Society for Information Science}, 49(11):973--982.

\bibitem[{Shojaee et~al.(2025)Shojaee, Mirzadeh, Alizadeh, Horton, Bengio, and Farajtabar}]{shojaee2025illusionthinkingunderstandingstrengths}
Parshin Shojaee, Iman Mirzadeh, Keivan Alizadeh, Maxwell Horton, Samy Bengio, and Mehrdad Farajtabar. 2025.
\newblock \href {https://arxiv.org/abs/2506.06941} {The illusion of thinking: Understanding the strengths and limitations of reasoning models via the lens of problem complexity}.
\newblock \emph{Preprint}, arXiv:2506.06941.

\bibitem[{Smith(2025)}]{smith2025sciencecomm}
Lucy Smith. 2025.
\newblock Science communication for {AI} researchers: Introductory training.
\newblock In \emph{Proceedings of the 39th AAAI Conference on Artificial Intelligence (AAAI‑25), Tutorial Track}.
\newblock 13:00–14:00 tutorial + 14:00–15:00 drop-in session.

\bibitem[{Srihari(1992)}]{156474}
S.N. Srihari. 1992.
\newblock \href {https://doi.org/10.1109/5.156474} {High-performance reading machines}.
\newblock \emph{Proceedings of the IEEE}, 80(7):1120--1132.

\bibitem[{Tait and Bensemann(2024)}]{tait2024clipping}
Izak Tait and Joshua Bensemann. 2024.
\newblock Clipping the risks: Integrating consciousness in {AGI} to avoid existential crises.
\newblock In \emph{International Conference on Artificial General Intelligence}, pages 176--182. Springer.

\bibitem[{Tweedie et~al.(1996)Tweedie, Singh, and Holmes}]{tweedie1996neural}
Fiona~J Tweedie, Sameer Singh, and David~I Holmes. 1996.
\newblock Neural network applications in stylometry: The federalist papers.
\newblock \emph{Computers and the Humanities}, 30:1--10.

\bibitem[{Waters(2025)}]{waters2025_diverging_future_ai}
Richard Waters. 2025.
\newblock The diverging future of {AI}.
\newblock \emph{Financial Times}.
\newblock Accessed: 2025-06-19.

\bibitem[{Weiser and Brown(1996)}]{weiser1996designing}
Mark Weiser and John~Seely Brown. 1996.
\newblock Designing calm technology.
\newblock \emph{PowerGrid Journal}, 1(1):75--85.

\bibitem[{Weizenbaum(1966)}]{weizenbaum1966eliza}
Joseph Weizenbaum. 1966.
\newblock Eliza—a computer program for the study of natural language communication between man and machine.
\newblock \emph{Communications of the ACM}, 9(1):36--45.

\bibitem[{Wyrich and Wagner(2023)}]{10.1109/ICSE-SEET58685.2023.00017}
Marvin Wyrich and Stefan Wagner. 2023.
\newblock \href {https://doi.org/10.1109/ICSE-SEET58685.2023.00017} {Teaching computer science students to communicate scientific findings more effectively}.
\newblock In \emph{Proceedings of the 45th International Conference on Software Engineering: Software Engineering Education and Training}, ICSE-SEET '23, page 107–114. IEEE Press.

\end{thebibliography}

\pdfoutput=1
\documentclass[11pt]{article}

\usepackage{etoolbox}
\newtoggle{anon}
\togglefalse{anon}

\iftoggle{anon}{\usepackage[review]{acl}}{\usepackage[final]{acl}}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{graphicx}

\title{NLP Meets the World: Toward Improving Conversations With the Public About Natural Language Processing Research}

\author{Shomir Wilson \\
  Department of Human-Centered Computing and Social Informatics \\
  Pennsylvania State University \\
  \texttt{shomir@psu.edu}\\}

\begin{document}
\maketitle
\begin{abstract}

Recent developments in large language models (LLMs) have been accompanied by rapidly growing public interest in natural language processing (NLP). This attention is reflected by major news venues, which sometimes invite NLP researchers to share their knowledge and views with a wide audience. Recognizing the opportunities of the present, for both the research field and for individual researchers, this paper shares recommendations for communicating with a general audience about the capabilities and limitations of NLP. These recommendations cover three themes: vague terminology as an obstacle to public understanding, unreasonable expectations as obstacles to sustainable growth, and ethical failures as obstacles to continued support. Published NLP research and popular news coverage are cited to illustrate these themes with examples. The recommendations promote effective, transparent communication with the general public about NLP, in order to strengthen public understanding and encourage support for research.

\end{abstract}

\section{Introduction}
\label{sec:intro}

Research communication has become a vital task for the natural language processing (NLP) research community, which faces a unique confluence of circumstances. Public interest in LLMs has led major news venues to publish non-technical explanations of how LLMs work~\cite{roose2023does, clarke2023_how_ai_chatbots_work}, predictions about the future of LLM development~\cite{waters2025_diverging_future_ai}, explorations of LLMs' sociodemographic biases~\cite{rogers2025_ai_bias_multilingual}, and concerns about their impacts on labor markets~\cite{hayes2025_ai_jobs}. News articles have also featured interviews with researchers who critique claims about artificial general intelligence (AGI)~\cite{heikkila2025_mitchell_agi_snake_oil, piquard2024_bender_parrots_lemonde}. Within scholarly literature, LLMs are sometimes claimed to be an existential threat to humanity~\cite{tait2024clipping, grozdanoff2023looming}, while others critique the plausibility or meaning of AGI~\cite{blilihamelin2025stoptreatingaginorthstar, mueller2024myth}. Regardless of stance, these topics and others reflect how NLP is experiencing an unprecedented surge of public attention.

The challenges of public discussion of NLP extend into culture and language. NLP researchers speaking with the public also contend with portrayals of AI in science fiction~\cite{osawa2022visions} and ambiguity over cognitive terms such as \textit{reason} and \textit{understand}~\cite{mitchell2023debate}. While research communication is a familiar task in any scholarly discipline, little support is available for the distinct set of challenges that NLP researchers currently face. This gap contrasts with the importance of research communication: it encourages public support for research funding, influences public policy, and engages the interest of students who may become future researchers.

This paper serves as a referenceable set of recommendations for effective conversations between NLP researchers and the general public. These recommendations are especially for researchers interacting with popular media (e.g., being interviewed for news articles) and promoting their work on social media. Three problems are described with accompanying guidance: \textit{vague terminology as an obstacle to public understanding} (\S\ref{sec:informative}), \textit{unreasonable expectations as obstacles to sustainable growth} (\S\ref{sec:learn}), and \textit{ethical failures as obstacles to continued support} (\S\ref{sec:specificity}). Rather than supplanting other guidance on research communication, this paper fills a notable gap by attending to challenges specific to the NLP research community.

\section{Related Work}

This paper follows an established practice of NLP conference papers that reflect on community challenges and suggest ways forward. Examples of this include \citeauthor{rogers-augenstein-2020-improve}'s recommendations for improving peer review of NLP papers~\citeyearpar{rogers-augenstein-2020-improve}, \citeauthor{blodgett-etal-2020-language}'s critique of how \textit{bias} is discussed by NLP researchers~\citeyearpar{blodgett-etal-2020-language}, and \citeauthor{mosbach-etal-2024-insights}'s discussion of interpretability research~\citeyearpar{mosbach-etal-2024-insights}. 

However, the published literature provides little guidance distinctly for NLP researchers on public communication (i.e., in contrast with guidance that applies to researchers in any discipline). Notably, \citet{hudson2023} observe that explainable AI researchers often rely on performance figures to describe their work, making it difficult to connect with a public audience. Conferences have hosted a series of tutorials on science communication for AI researchers, most recently at AAAI 2025~\cite{smith2025sciencecomm}, but the reach and accessibility of those tutorials differ from a published paper. \citet{10.1109/ICSE-SEET58685.2023.00017} designed a course-based exercise for computer science undergraduates to practice communicating the findings of a paper from software engineering, a related field. However, they focus on describing the exercise itself rather than sharing guidance on its topic.

In the broader realm of research communication, several works provide guidance that is generally applicable to NLP but do not address the challenges that NLP researchers currently face. \citet{fontaine2019communicating} surveyed the literature on public communication by researchers within the health sciences. They noted wide use of social media and proposed a typology of strategies. These include minimizing the use of jargon, encouraging public discussion, and making information actionable, among others. The US National Academies \citeyearpar{national2017communicating} provided extensive guidance on science communication topics such as handling controversy, understanding one's audience, and using social media effectively. \citet{kuncel2013communicating} address science communication for psychology, with a focus on making statistical methods understandable.

All of the above works are valuable contributions, and it is appropriate to select among them for a basic understanding of research communication. This manuscript builds upon them by focusing on the unique circumstances of NLP research, remedying a lack of field-specific guidance.

\section{Vague Terminology as an Obstacle to Public Understanding}
\label{sec:informative}

NLP researchers often describe their work using terms that public audiences associate with cognition. Sometimes researchers do not intend to make cognitive claims with these terms, but sometimes the cognitive claims are intentional. Additionally, the research community lacks an agreed-upon stance on some cognitive claims, partly because of a lack of agreement for what they mean~\cite{bender-koller-2020-climbing}. These conditions are difficult for non-experts to navigate. This section illustrates with terms used in published research.

\subsection{It's Complicated: NLP's Relationship With Cognitive Terms}

Terms with cognitive implications (among their other implications) have a long history in NLP, to the extent that some are widely interpreted by researchers without referring to cognition. \textit{Predict}, for example, has been associated with text classification at least as early as the 1990s~\cite{riloff1992classifying, 10.1145/183422.183425} and possibly earlier. \textit{Read} also has a long history~\cite{156474, GOVINDAN1990671}. \textit{Learn} (as in \textit{machine learning}) originates from a 1959 paper by Arthur Samuel about programs that play the game of checkers~\cite{10.1147/rd.33.0210, foote2021briefml}. However, a non-expert may associate all three of those terms with the ability to \textit{think} or \textit{understand}, broadly without definitions. Observe that ``a human must understand to predict'' is intuitively reasonable, but within the NLP research community ``a language model must understand to predict'' is controversial within the framing of several papers cited in \S\ref{sec:intro}.

Recently LLMs' capabilities and limitations have amplified attention to cognitive terms. \textit{Hallucinate} is widely recognized by the research community~\cite{huang2025survey}, even if a precise meaning is not agreed upon~\cite{narayanan-venkit-etal-2024-audit}. However, the relevance of \textit{understand}, \textit{reason}, and \textit{think} are controversial~\cite{bender-koller-2020-climbing, shojaee2025illusionthinkingunderstandingstrengths}. Again, the relationships between these terms are complex within NLP, to an extent unknowable by an outside audience. A human who is \textit{hallucinating} is presumably \textit{thinking} (albeit aberratively), but a \textit{hallucinating} LLM is not agreed upon by the research community to be \textit{thinking}. A further complication happens when arguments against LLMs' cognitive abilities nominally exclude humans from having those same abilities. For example, \citet{shojaee2025illusionthinkingunderstandingstrengths} argue that language reasoning models (LRMs, closely related to LLMs) cannot \textit{think} because of ``a complete accuracy collapse'' for complex tasks. The human brain also has task complexity limits~\cite{marois2005capacity}, but in contrast with LLMs, it is widely recognized as capable of \textit{thinking}.

\subsection{Recommendations}

NLP researchers speaking with the public should provide scope for cognitive terms they use. Given the above obstacles, it is helpful to tell a general audience what terms such as \textit{predict} or \textit{reason} mean in a context. If a non-cognitive term can be used instead of a cognitive one, that substitution may provide greater clarity and transparency. When benchmarks are relevant, describing their role can demystify how a cognitive or pseudo-cognitive claim is justified. A general goal is to avoid the possibility of unintended cognitive implications for an LLM or other human language technology.

NLP researchers should balance correcting cognitive misconceptions about LLMs with tolerating a general audience's verbal shorthand and metaphor. Humans apply cognitive anthropomorphism to a variety of objects~\cite{boyer1996makes} such as lamps and umbrellas~\cite{edwards2022lamps}, and LLMs display far more sophisticated behavior than those objects. When cognitive terms surface in discussions about LLMs, a diplomatic strategy is to point out the verbal shorthand or metaphor, acknowledge its naturalness, and explain the actual limitations of the technology. This strategy validates a non-expert's experiences with technology while illuminating how cognitive terms can mislead.

\section{Unreasonable Expectations as Obstacles to Sustainable Growth}
\label{sec:learn}

Predictions of the impacts of LLMs include radical improvements in biology, neuroscience, economic development, and human welfare, and the onset of global peace~\cite{amodei2024machines}. The possibility of an ``AI bubble''\footnote{There exists a lack of agreement on the distinction between AI and NLP, with some common ground and some differences. To avoid misattribution, this paper uses ``AI'' for topics where the term currently prevails in popular discussion.} (i.e., an unsustainable level of public interest in AI that exceeds its capabilities) is widely acknowledged~\cite{horowitch2025computersciencebubble, press2025peakai}. A bubble implies short-term benefits, such as publicity and abundant resources. However, its collapse---implied by the metaphor---creates harm through an unpredictably sharp decline in resources. This section provides examples from the history of AI and recommendations for how NLP researchers can encourage sustainable perceptions of the field.

\subsection{Remembering AI Winters}

Multiple AI winters are said to have occurred, and they lack agreed-upon dates. Roger Schank and Marvin Minsky referred to one in the 1970s~\cite{crevier1993ai}, and James Hendler referred to one in the 1980s~\cite{4475849}. This section uses Hendler's perspective to structure the discussion.

Hendler described a confluence of three conditions that led to the AI winter of the 1980s. The first condition was a decrease in government funding for AI research in the 1970s, specifically in the USA, which dominated AI research at the time. He argues the impact was delayed by five to ten years, as already-approved projects were permitted to run their course. For the rationale of the cuts, Hendler quotes a 1973 report by James Lighthill in the UK that concluded ``in no part of the field have discoveries made so far produced the major impact that was then promised.''~\cite{james1973artificial} The second condition was an erosion of recognition for AI's successes. Hendler argues that expert systems~\cite{buchanan1988fundamentals}, which were successful in industry, were ``disowned'' by AI researchers in favor of topics where less progress had been made. The third condition was the rapid democratization of hardware needed to run AI models, in the sense that the hardware became dramatically cheaper and easier for individuals to own, thus disrupting sociotechnical power structures within research~\cite{4475849}.

These conditions require some subjectivity to assess, making it difficult to firmly state whether they exist now or will appear in the near future. However, there are discernible lessons in them for how NLP researchers communicate their work.

\subsection{Recommendations}

For the NLP research community as a whole, the downsides of moderating expectations are modest when compared to the downsides of a sudden collapse in public interest. At the individual level, one researcher's careful moderation when speaking about their work is unlikely to change public perception of the field, but strategies exist that fulfill that moderation while creating individual benefit as well. Such a strategy is described below.

Public engagement is an opportunity to support recognition of NLP applications that may have been inadvertently ``disowned'' (borrowing Hendler's description) for their success. NLP played a formative role in early search engines~\cite{schwartz1998web} and enabled early spam detection~\cite{cormack2008email}. Optical character recognition made a wealth of pre-digital texts searchable~\cite{hill2019quantifying}, and authorship identification solved historical mysteries~\cite{tweedie1996neural}. Those examples and others show NLP applications as \textit{calm technologies}, supporting human work rather than monopolizing human attention~\cite{weiser1996designing}. They also may serve as an antidote to hype by illustrating the expectation that research products need time to mature. Dialog systems are a strong public-facing example of the long road to practical use: early systems such as Eliza~\cite{weizenbaum1966eliza} are distant ancestors of modern virtual assistants, such as Amazon Alexa and Google Assistant.

\section{Ethical Failures as Obstacles to Continued Support}
\label{sec:specificity}
Ethically dubious practices are known to exist within the NLP research community, and outside of it, problematic applications exist as well. Both are obstacles to how NLP research is received by a public audience. This section gives examples of relevant ethical failures and provides suggestions for how NLP researchers can speak frankly and constructively when called upon to discuss them with the public.

\subsection{Reckoning With Problematic NLP}

Research institutions in many countries have procedures for reviewing experimental protocols that involve human subjects. However, sometimes an experiment complies with requirements while violating ethical norms, or researchers fail to seek approval. A recent example is an experiment in which researchers studied the persuasiveness of an LLM-powered chatbot by covertly deploying it on the Reddit forum \textit{r/changemyview}\footnote{https://www.reddit.com/r/changemyview/}, concealing that it was an LLM. When forum moderators were told of the deception, they expressed shock, asked for an apology from the university, and requested that the research not be published~\cite{OGrady2025_unethical_AI_Reddit}.

A comprehensive list of ethical failures or misuses of NLP applications is impractical, but examples are readily available. LLMs have been used to generate disinformation, potentially manipulating geopolitics~\cite{Fried2025}. Chatbots have demonstrated the capacity to respond harmfully in conversations with people suffering from mental illness~\cite{coghlan2023chat}, and more broadly, LLMs' sociodemographic biases have been shown to harm individuals from minoritized groups~\cite{Ghosh2024}. The broad-based collection of training data for LLMs raises privacy concerns and intellectual property concerns~\cite{novelli2024generative}. LLM-assisted writing has been shown to produce ``cognitive debt'' and reduce critical thinking~\cite{kosmyna2025brainchatgptaccumulationcognitive}. 

The above harms and others receive public attention, to the extent that a 2023 Pew Research survey found 52\% of Americans were more concerned than excited about AI in their daily lives~\cite{FaverioTyson2023}. This level of concern is demonstrably relevant to NLP researchers.

\subsection{Recommendations}

Several popular news articles cited in the above section show that NLP researchers are called upon to discuss problematic uses of NLP. Open dialog about these problems is obligatory and familiar as part of open science, but further steps can help to sustain public support. Researchers speaking with the general public about problematic NLP should be prepared to explain what should have been done differently to achieve the same goals, or why the goals should never have been undertaken. Public interaction is also an opportunity to point to research efforts directly aimed at ethical issues in NLP, showing how human-centered research is a valuable part of the community. Such research is a marked contrast with a hypothetical public image of AI as solely focused on computing.

Within the research community, researchers can lay the groundwork for strong public support by recognizing that research that critiques NLP fully belongs in the community. Such research does not necessarily create human language technologies, but it guides their ethical development. This inclusion also encourages grounding NLP research in human benefit, which demonstrates responsiveness to the concerns of the general public.

\section{Conclusion}

Public attention to NLP research comes with benefits, obstacles, and responsibilities. This paper presented recommendations to help NLP researchers communicate effectively with the public, focusing on obstacles that are distinctive to our research community. It drew upon a variety of sources from research literature and popular news coverage of NLP to illustrate those obstacles and how to navigate them, for the prosperity of both individual researchers and the research community at large.

\iftoggle{anon}{}{

\section*{Acknowledgments}

Kenneth Huang and Natalie Parde are gratefully acknowledged for their feedback on ideas in this manuscript. This manuscript is based upon work supported by the National Science Foundation under Grant \#2237574.

} % close anon iftoggle

\section*{Limitations}

This manuscript is not a comprehensive guide to research communication, and instead it augments such guides. In \S1 and \S2 we remind readers that our contribution is an NLP-specific addition to existing literature about research communication, and in \S2 we cite several sources the reader may consult for more background.

NLP researchers also may face public communication challenges not covered in this guide. We identify three challenges of particular concern, but others are likely to exist. The subjective nature of the topic space makes it difficult to segment it in a principled way, but we suggest that providing evidence-based guidance (as in this manuscript) in the published literature is more likely to lead to positive outcomes than a lack of guidance.

%Since December 2023, a "Limitations" section has been required for all papers submitted to ACL Rolling Review (ARR). This section should be placed at the end of the paper, before the references. The "Limitations" section (along with, optionally, a section for ethical considerations) may be up to one page and will not count toward the final page limit. Note that these files may be used by venues that do not rely on ARR so it is recommended to verify the requirement of a "Limitations" section and other criteria with the venue in question.

% Bibliography entries for the entire Anthology, followed by custom entries
%\bibliography{anthology,custom}
% Custom bibliography entries only
\bibliography{aclrr2025_shomir}

\end{document}